I am struggling to fufil my teacher's task. She asked me to right out all the steps to slove the following problems using the quadratic formula. The problems are x^2+20x+100=64 , 4x^2+4x+1=16, 36x^2+12x+1=18, 9x^2-12x+4=4, and 25x^2 - 30x+9=96. I know this is alot. I am seriously struggling in Algebra 2 and I am in great need of a tutor.
Answer by edjones(8007)

x^2+20x+100=64
x^2+20x+36=0
(x+18)(x+2)=0
x=-18, x=-2
.
Ed